a rocket with a mass of 400 lb travels at 27000 ft per sec. what is its kinetic energy in ft-lbf?
1
Expert's answer
2019-06-19T10:29:49-0400

The kinetic energy:


"K=0.5mv^2""m=400\ lb_m=\frac{400}{32.174}slugs"


"K=0.5\frac{400}{32.174}(27000)^2=4.5\cdot 10^9\ ft \cdot lb_f"